Records over the past year show that 1 out of 355 loans made by Mammon Bank have defaulted. Find the probability that 2 or more out of 250 loans will default. Hint: Is it appropriate to use the Poisson approximation to the binomial distribution? (Round λ to 1 decimal place. Use 4 decimal places for your answer.)
Answer #1

here expected number of people who will default in 250 loans =250/355=0.7

hence probability that 2 or more out of 250 loans will default =P(X>=2)=1-P(X<=1)

=1-(P(X=0)+P(X=1))=1-(e-0.7*0.70/0!+e-0.7*0.71/1!)=0.1558